80% of 35
16% of 70
150% of 80
118% of 3,000
5% of 12
1
Expert's answer
2011-05-17T08:08:02-0400
35*80%/100% = 28
70 * 16% / 100% = 11.2
80 * 150%/ 100% = 120
3000 * 118% /100% = 3540
12 * 5%/100% = 0.6